OpenStudy (anonymous):

Solve. 4y = 144 36 16 –16 –36

OpenStudy (anonymous):

@kaedeedee @DDcamp

OpenStudy (anonymous):

@kadeedee

OpenStudy (texaschic101):

divide both sides by 4 and what do you get ?

OpenStudy (anonymous):

y=36

OpenStudy (texaschic101):

4y = 144 4y/4 = 144/4 y = 36